THE QUINTET OF HF DIMER: VIBRATIONAL OVERTONES OF THE HF STRETCHES AT N = 4

\underline{ZHENHONG YU}, ESSAM HAMMAM, and WILLIAM KLEMPERER, Harvard University, Department of Chemistry and Chemical Biology, Cambridge, Massachusetts 02138.

The vibrational overtones of HF dimer at N = v_{1} + v_{2} = 4 is a quintet: (4, 0), (3, 1), (2, 2), (1, 3), and (0, 4), in the visible region. The K = 0 and 1 sublevels of the (4, 0) mode has been observed near 14,700 cm^{−1} by the high-resolution molecular beam intra-cavity laser induced fluorescence. The band origin ν_{0} = 14700.458(7) cm^{−1} and rotational constant (B + C)/2 = 0.22278(31) cm^{−1} are determined for K = 0 with the centrifugal distorsion constant D_{J} fixed to 2 × 10^{−6} cm^{−1}. The spectral constants A = 24.3 cm^{−1}, (B + C)/2 = 0.22296(20) cm^{−1}, and (B - C) = 4.5(2) × 10^{−3} cm^{−1} are also obtained for K = 1. The hydrogen interchange tunneling is essentially quenched at both K = 0 and 1 levels, consistent with the theoretical prediction from a phenomenological model.^{a} The predissociation linewidths are 470(30) MHz with no apparent rotational dependence for both levels. To further verify the integrity of our model, we are extending our experimental studies to the (3, 1) and (2, 2) modes.

aH. -C. Chang and William Klemperer, J. Chem. Phys. 104, 7830(1996).